BREAKDOWN OF THE REDUCTION OF THE ROVIBRATIONAL HAMILTONIAN : THE CASE OF S^{18}O_{2}F_{2}

\underline{L. MARGULÈS}, J. DEMAISON, Laboratoire PhLAM, CNRS UMR 8523, Université de Lille 1, Bat. P5, 59655 Villeneuve d’Ascq Cedex, France.; A. PERRIN, Laboratoire Inter Universitaire des Systemes Atmosphériques, CNRS UMR 7583, Université Paris 12, 61 Av du General de Gaulle, 94010 Créteil Cedex France.; I. MERKE, ; Institute für Physikalische Chemie, RWTH Aachen, 52056 Aachen, Germany; H. WILLNER, Anorganische Chemie, FB C, Universität-GH Wuppertal, 42097 Wuppertal, Germany; M. ROTGER, Groupe de Spectromtrie Moléculaire et Atmosphrique, CNRS UMR 6089, Moulin de la Housse, BP 1039, Cases 16-17, 51687 Reims Cedex 2, France; and V. BOUDON, Institut Carnot de Bourgogne, UMR CNRS 5209, 9 avenue Alain Savary, BP 47870, 21078 Dijon Cedex, France.

The ground state rotational spectrum of the near-spherical top molecule S^{18}O_{2}F_{2} (sulfuryl fluoride) has been measured from 50 to 700 GHz. As for the parent isotopologue, S^{16}O_{2}F_{2}^{a}, it was necessary to use a non-reduced Hamiltonian in order to obtain a satisfactory fit. It was possible to determine six quartic centrifugal distortion constants (instead of five for a standard asymmetric top) and an additional sextic constant could also be determined. This ground state level has also been analysed thanks to a tensorial formalism developed in Dijon. Only two tensorial sextic constants are fixed to zero, all others have been adjusted. Although S^{18}O_{2}F_{2} is less spherical than S^{16}O_{2}F_{2}, the analysis was more difficult. It is partly due to the fact that S^{18}O_{2}F_{2} is oblate whereas S^{16}O_{2}F_{2} is prolate. The experimental quartic centrifugal distortion constants were found in good agreement with those calculated from the force field, confirming the correctness of the analysis.

aK. Sarka, J. Demaison, L. Margulès, I. Merke, N. Heineking, H. Bürger, and H. Ruland, J. Mol. Spectrosc., 200, 55 (2000)